what is the perimeter when you have a length of 9 and a area of 126 square feet
Answer by macston(5194)

A=area=126 sq ft; P=perimeter; L=length=9ft; W=width
A=L*W
A/L=W
126sqft/9ft=W
14 ft=W
P=2(L+W)=2(9ft+14ft)=2(23ft)=46 feet ANSWER: The perimeter is 46 feet.